PLEASE HELP ME SOLVE THE FOLLOWING PROBLEM. IF YOUR TOTAL INVOICE IS 17.42 WHAT WAS THE SUBTOTAL IF THE SALES TAX WAS 7%
Found 2 solutions by b_howie, stanbon:
Answer by b_howie(8)

(x/17.42)=(93/100)
100x=93(17.42)
100x=1620.06
16.20=x
Subtotal:$16.20

Answer by stanbon(57361)

IF YOUR TOTAL INVOICE IS 17.42 WHAT WAS THE SUBTOTAL IF THE SALES TAX WAS 7%
Let the sales price be "x".
x+0.07x=17.42
1.07x=17.42
x=$16.28
============
Cheers,
Stan H.